INFRARED RESONANCE ENHANCED PHOTODISSOCIATION SPECTROSCOPY OF Si^{+}(CO_{2})_{n}, Si^{+}(CO_{2})_{n}Ar, AND Si^{+}(O_{2})(CO_{2})_{n} CLUSTERS

\underline{J.B. JAEGER}, T.D. JAEGER, and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

Weakly bound Si^{+}(CO_{2})_{n}, Si^{+}(CO_{2})_{n}Ar, and Si^{+}(O_{2})(CO_{2})_{n} complexes are produced by laser vaporization in a pulsed supersonic expansion. The ions are then mass selected in a reflectron time-of-flight mass spectrometer where they are photodissociated by a tunable infrared OPO/OPA laser system near the asymmetric stretch (ν_{3}) of CO_{2}. Dissociation is more efficient on resonance, thus monitoring the fragmentation channel as a function of infrared laser frequency produces the IR absorption spectrum for the complex of interest. Si^{+}(CO_{2})_{n} and Si^{+}(O_{2})(CO_{2})_{n} clusters fragment by the loss of whole CO_{2} units while Si^{+}(CO_{2})_{n}Ar complexes fragment by the loss of argon. Interpretation of the observed infrared spectra will be discussed.